OpenStudy (anonymous):

8y - 9 = 7/2y ((show work))

OpenStudy (anonymous):

2(8y - 9) = (7/2)(y)(2) Multiplying both sides by 2 16y - 18 = 7y Simplifying 9y - 18 = 0 Subtract both sides by 9y 9y = 18 Add 18 to both sides y = 2 Divide both sides by 9